the efficiency of a machine is 60%. find in joules the workdone by a student using this machine to raise a load of 150kg through a vertical distance of 2.5m
1
Expert's answer
2018-05-15T15:31:07-0400
A_Useful=F·s=m·g·h
A_efficiency=A_Useful/ɳ=(m·g·h)/ɳ=(150 ·9.81·2.5)/0.6=6131.25 Joules